(b): Finding probabilities for normal distribution

i. Given a population with a normal distribution, a mean of 0, and a standard deviation of 1, find the probability of a value less than 0.75.

ii. Given a population with a normal distribution, a mean of 120, and a standard deviation of 10, find the probability of a value greater than 135.

iii. Given a population with a normal distribution, a mean of 98.6, and a standard deviation of 0.62, find the probability of a value between 97.0 and 99.0.

iv. Given a population with a normal distribution, a mean of u = 17 and the standard deviation of 3, find the 85th percentile.


1
Expert's answer
2021-08-11T08:07:12-0400

(b)

i.


"P(Z<0.75)\approx0.773373"

ii.


"P(X>135)=1-P(X\leq 135)"

"=1-P(Z\leq\dfrac{135-120}{10})=1-P(Z\leq1.5)"

"\approx0.066807"

iii.


"P(97.0<X<99.0)=P(X<99.0)-P(X\leq 97.0)"

"=P(Z<\dfrac{99.0-98.6}{0.62})-P(Z\leq\dfrac{97.0-98.6}{0.62})"

"\approx P(Z<0.6451613)-P(Z\leq-2.5806452)"

"\approx0.7405889-0.0049308"

"\approx0.735658"

iv.


"P(Z<z)=0.85"

"z\approx1.036433"

"\dfrac{x-u}{\sigma}=1.036433"

"x=17+3(1.036433)"


"z\approx20.1093"

The 85th percentile is "20.1093."